# Advancing Large Multi-modal Models with Explicit Chain-of-Reasoning and Visual Question Generation

### v1

@@ Abstract
The increasing demand for intelligent systems capable of interpreting and reasoning about visual content requires the development of Large Multi-Modal Models (LMMs) that are not only accurate but also have explicit reasoning capabilities. This paper presents a novel approach to imbue an LMM with the ability to conduct explicit reasoning based on visual content and textual instructions. We introduce a system that can ask a question to acquire necessary knowledge, thereby enhancing the robustness and explicability of the reasoning process. Our method comprises the development of a novel dataset generated by a Large Language Model (LLM), designed to promote chain-of-thought reasoning combined with a question-asking mechanism. We designed an LMM, which has high capabilities on region awareness to address the intricate requirements of image-text alignment. The model undergoes a three-stage training phase, starting with large-scale image-text alignment using a large-scale datasets, followed by instruction tuning, and fine-tuning with a focus on chain-of-thought reasoning. The results demonstrate a stride toward a more robust, accurate, and interpretable LMM, capable of reasoning explicitly and seeking information proactively when confronted with ambiguous visual input.
@@ 1 Introduction
In recent years, Large Language Models (LLMs) have gained significant popularity in the field of artificial intelligence [5, 35, 48]. Building upon this, research has progressed into Large-scale Multi-modal Models (LMMs) [21, 11, 27, 26]. Our study specifically focuses on tasks involving both vision and language modalities. Generally, LMMs learn the alignment between images and texts using large datasets sourced by web crawling. They are then fine-tuned with smaller, high-quality datasets for instruction tuning to enhance their text generation capabilities for various tasks. However, these models often struggle with hallucination, where they produce outputs not aligned with the given input. For instance, in LMMs, the models may mention objects that do not exist in the given images [25, 24].
@@
Therefore, our research aims to address these challenges by incorporating an explicit reasoning process and the ability to generate questions during reasoning. This approach is akin to the Chain-of-Thought reasoning seen in LLMs [52, 10], where models are prompted to explicitly reason out problems. However, due to LMMs’ relatively weaker text generation capabilities compared to LLMs, achieving Chain-of-Thought solely through prompting is challenging. To overcome this, we have created a novel dataset that includes the explicit reasoning process. Following mainstream methods in LLM-based dataset creation [27], we generate this data using LLMs, combining image annotations and a few manually created examples. This data includes scenarios where the model needs to generate questions when uncertain, training it to ask questions during ambiguous reasoning situations. We refer to this reasoning process as “Chain-of-Reasoning (CoR).”
We designed a novel LMM architecture to be trained on this dataset, which has several unique features, such as improved region-awareness and integrating pre-trained image encoder and text decoder. Being trained on our dataset, the model acquires the ability to generate explicit reasoning steps and ask questions when uncertain, thereby improving the reliability of its inferences. Our contributions are summarized as follows:
We present an innovative approach by incorporating an explicit reasoning process and question-generation capability into LMMs, promoting more reliable inferences.
We crafted a new dataset and utilized it for model training, setting a precedent for future LMM advancements.
@@ 2 Related Work
2.1 Large-scale Multi-modal Models
@@
BLIP-2 [21] and InstructBLIP [11] utilize pre-trained image encoders and text decoders, employing a Transformer model known as Q-Former as an adapter. The training of these models is typically divided into two stages. In the first stage, Q-Former learns the alignment between image and text using large-scale image-text paired datasets. The second stage involves training the Q-Former alongside the text decoder to enable the decoder to produce outputs according to given instructions. Generally, while the first training stage prioritizes the scale of the dataset, the second stage emphasizes the quality of smaller, high-quality instruction and answer datasets. This two-stage learning approach is widely adopted in the training of LMMs, as seen in models like BLIVA [17], MiniGPT (v1&2) [58, 8], and LLaVA (v1&1.5) [27, 26].
Additionally, there’s growing interest in LMMs that not only produce outputs aligned with overall image instructions but also focus on specific regions within images. This research direction is evident in models such as KOSMOS-2 [38], Shikra [9], MiniGPT v2 [8], and LLaVA v1.5 [26]. These models transform bounding box coordinates, representing input regions, into special tokens for textual representation. Another noteworthy model is GPT4RoI [55], which applies an RoIAlign based on the target region to the features outputted by the image encoder, thus obtaining features focused on the specified region.
@@ 2.3 Visual Question Generation
A related field to the concept of acquiring information through question generation is Visual Question Generation (VQG). In the early stages of VQG, the primary focus was on generating questions related to images, without any specific goal of acquiring knowledge [33]. K-VQG [49] does involve generating questions for knowledge acquisition; however, there has been no emphasis on utilizing the acquired knowledge in further reasoning steps.
Learning by Asking (LBA) is another relevant area. LBA is an approach where question generation is used to acquire data for learning purposes. For instance, LBA has been employed in tasks like VQA [32], caption generation [44], and scene graph completion [53]. However, our research significantly diverges from LBA, as we focus on generating questions as part of explicit reasoning during the inference process. The acquired answers are then immediately used to update and refine the reasoning quality, differentiating our approach from traditional LBA work.
@@ 3.1 Dataset Construction
In pursuit of realizing the CoR, we developed a novel dataset. This dataset was created using a combination of the specially designed prompt, manually curated examples and the text generation capabilities of LLM. We show an example of the dataset in Figure 2.
The data structure of our dataset consists of the image, an instruction text of the task, a question (optional, if the task is in QA format), the answer, and the reasoning process required to derive the answer. Given that LLMs do not inherently accept image inputs, we employ text-based format to encapsulate the content of the image, following LLaVA’s approach [27]. This is achieved by providing the LLM with annotations of the image, which include captions (around five sentences), object bounding boxes, and their corresponding labels. These textual inputs serve as a proxy, enabling the LLM to understand and process image content without direct image input. We show an overview of the dataset construction process in Figure 3.
@@
Coupling these inputs with thoughtfully designed prompts enables the LLM to generate the appropriate reasoning steps. It’s crucial that the prompts ensure that the LLM simulates the act of observing the image directly, without mentioning any references to captions, descriptions, or bounding boxes. This approach ensures the consistency of the reasoning process when training LMMs, which do view the image directly. To achieve this, in the prompt given to the LLM, we clearly state like you must not use the word “caption”, “description”, or “bounding box” in the reasoning steps. Especially, do not mention the coordinates of each object in the image. If you want to mention the position of the object, please refer to it by relative position, such as “at the left of the image” or “in front of xx”. Furthermore, in the prompt, we instruct the LLM to suppose that the image is of size 10001000 pixels. This ensures accurate position estimations based on given textual coordinates of the bounding boxes. The prompts also emphasize succinct reasoning steps without redundancy, and the need to add uncertainty scores at the end of each reasoning step. We show the full prompt in Listing LABEL:lst:prompt.
@@
We devised three variants of reasoning step data: without QA, with QA, and with GT. In “without QA” setting, the reasoning process to the answer is conducted without generating any questions midway. This setting aims to mimic the process an LMM undergoes when attempting to answer a question directly, without generating any questions. In this scenario, the LLM is instructed to generate reasoning steps without seeing the correct answer to the given task. We instruct the model to generate “Uncertainty score” along with each reasoning step, indicating the model’s confidence in its reasoning. This plays a crucial role in “with QA” setting described below.
@@
In contrast, the “with QA” setting incorporates the generation of questions as part of the reasoning steps. This setup is crucial for training LMMs to perform explicit reasoning while simultaneously generating questions. As the model generates the reasoning steps, it also generates an uncertainty score for each step. If the score significantly rises at any point in the “without QA” setting, it triggers the generation of a question just before the step. This mechanism helps the model to identify moments of uncertainty and seek external information, thereby simulating a more interactive and dynamic reasoning process.
@@
Lastly, the “with GT” setting involves the LLM generating reasoning steps while having access to the correct answers to the tasks (i.e., the ground-truth answers for the given questions or captions for the given images). This approach differs from the previous two as the reasoning steps invariably lead to the correct answer, serving as a ground truth for correct reasoning without question generation. This data is primarily used to train LMMs that do not involve question generation in their reasoning process.
@@
COCO Caption | 5,938 | 5,863 | 6.27 | 8.88 | 6.20 |
VQA v2 | 5,948 | 5,813 | 4.69 | 7.61 | 4.55 |
OK-VQA | 5,941 | 5,940 | 4.77 | 7.58 | 4.55 |
A-OKVQA | 5,946 | 5,927 | 4.87 | 7.66 | 4.63 |
Visual Genome Caption | 5,973 | 5,696 | 4.34 | 7.31 | 4.12 |
Total | 29,746 | 28,042 | 4.99 | 7.81 | 4.81 |
@@ 3.2 Dataset Statistics
We provide a comprehensive overview of our dataset statistics in Table 1. In our research, we based our dataset generation on five datasets: COCO Caption, VQA v2, OK-VQA, A-OKVQA, and Visual Genome Caption. For each of these datasets, we generated approximately 6,000 data samples, randomly selecting from their training sets.
When generating reasoning data without questions and answers (without QA) and when not seeing ground-truth answers (with GT), the average number of reasoning steps is shorter compared to the scenario where both questions and answers are included (with QA). This reduction in steps is attributed to the fact that the QA generation process inherently requires more reasoning steps.
@@
Furthermore, the COCO Caption dataset tends to have longer reasoning steps compared to the others. This is possibly because the COCO Caption dataset involve generating captions based on the entire content of the image, requiring a more comprehensive reasoning process. In contrast, the other tasks focus on specific questions or target regions within the images, thereby requiring less information to be incorporated into the reasoning steps.
@@ 3.3 Model Architecture
Our model fundamentally builds upon the BLIP-2 architecture [21], consisting of an image encoder, a text decoder, and an adapter component termed the Q-Former. We show an overview of our model in Figure 4.
Image Encoder: We use a pre-trained CLIP-ViT [39] to extract image features. A novel aspect we introduced is the dual-input system: the model not only receives the entire image but also another version where the image except the region of interest are masked. This strategic input combination allows for more region-aware text generation compared to conventional methods.
Q-Former: Serving as the bridge between the image encoder and text decoder, the Q-Former takes the image features and converts them to a feature suitable for the text decoder. Its underlying structure is based on the Transformer architecture. The Q-Former processes both the image features and a set of trainable query tokens. Its training is governed by three distinct loss functions: the text-image matching loss, the text-image contrastive loss, and the text generation loss. Please refer to the original BLIP-2 paper [21] for more details on the Q-Former.
Text Decoder: The text decoder is designed to receive features from both the Q-Former and linearly projected image features. By accepting raw features from the image encoder in addition to the Q-Former’s outputs, the decoder benefits from a richer information about the iamge. This design choice was inspired by the BLIVA model [17], and we adopted their proposed method to ensure a more comprehensive input for the decoder.
4 Training
Our training process consists of three stages: (1) image-text alignment, (2) instruction tuning, and (3) CoR fine-tuning.
4.1 Stage1: Large-scale Image-text Alignment
The first stage of training is large-scale image-text alignment. In this phase, we utilized a mixture of web-curated and manually annotated datasets listed as follows:
- MS COCO
[22] is a large-scale dataset consisting of about 80K images with 5 captions per image.
- Visual Genome
- RECO
[54] is another region-aware dense caption dataset. This dataset is created for region-controlled text-to-image generation task. The images are from MS COCO and LVIS [13], and the captions are automatically generated by GIT [51]. Same as Visual Genome, we remove captions associated with small region.
- GRIT
[38] is also a region-aware dense caption dataset. This dataset is based on COYO-700M [6] and LAION-2B [41] dataset. The original captions in them are automatically refined by detection model and post-processing pipeline to produce region-aware captions in massive scale. We used published version of GRIT, which is a subset of the original dataset and contains samples from COYO-700M.
- SBU captions
[36] is a image-caption dataset consisting of about 850K images and captions.
- Conceptual Captions 3M
[43] is a dataset consisting of about 3M images and captions. The data are collected from the web-crawled images and their associated alt texts, and applied with a filtering process to remove unsuitable data.
- LAION 400M
For region-aware image-text datasets, we used the bounding box annotations to mask the image except the region of interest. In this stage, we froze the parameters of the image encoder, and only trained the Q-Former.
4.2 Stage2: Instruction Tuning
In Stage 2, we focus on refining the model’s ability to perform tasks based on a wide array of instructional texts through instruction tuning. Here, it is crucial to train the model with datasets that include a diversity of instructions and tasks to enhance its adaptability and understanding. For this phase, we employed an assortment of different datasets to instruct the model across a spectrum of tasks. The datasets utilized and the corresponding tasks they encompass are listed as follows:
Caption Generation: This task involves generating captions to images, utilizing datasets like MS COCO [22], SVIT [57], and VizWiz-caption [16]. MS COCO tasks the model with generating captions with single sentence, SVIT requires longer captions comprising several sentences, and VizWiz-caption focuses on generating captions for visually impaired individuals.
Region Caption Generation: In this task, models generate captions for specific image regions. Datasets used include Visual Genome [20] and RefCOCO [19, 30], which are manually curated and typically feature short captions of a few words. Meanwhile, GRIT [38] and Reco [54] represent datasets where captions have been auto-generated by existing caption generation models.
Image Comprehension: VSR [23], IconQA [28], and ScienceQA [29] are the datasets chosen for this task category. VSR involves determining the veracity of statements describing the contents of an image, while IconQA and ScienceQA require the model to answer questions about images by choosing from multiple options.
Knowledge-Based VQA: For tasks like OK-VQA [31] and A-OKVQA [42], the model must tap into external knowledge sources to generate appropriate answers.
VQA: Standard VQA tasks employ datasets like VQA v2 [12], GQA [18], and VizWiz-VQA [15]. GQA is notable for its automatically generated questions based on scene graphs, many of which involve spatial reasoning. VizWiz-VQA poses a unique challenge as it contains questions from visually impaired users, requiring text reading in images or dealing with poor image quality.
OCR Tasks: TextVQA [46] and TextCaps [45] require the model to read and interpret text within images for VQA or caption generation.
Moreover, we integrated recently proposed datasets specifically designed for Instruction Tuning -— Polite Flamingo [7], LRV Instructions [25], LLaVA [27], and LLaVAR [56] -— into our training regime. In this stage, we froze the parameters of the image encoder and text decoder, and only updated the parameters of the Q-Former. The model is trained to minimize the cross-entropy loss between the predicted tokens and the ground-truth tokens.
4.3 Stage3: CoR Fine-tuning
In stage 3, we utilizes our novel dataset to make the model learn to generate reasoning steps and questions. This stage is furthur divided into two substeps:
Substep 3-1. In this phase, the model learns to generate reasoning steps and questions. This phase is dedicated to teaching the model how to craft reasoning steps that naturally lead to the generation of questions. It targets the ability to formulate queries as part of the reasoning process, an essential step for interactive problem-solving.
MS COCO | Visual Genome | OK-VQA | A-OKVQA | |||||
BLEU-4 | METEOR | CIDEr | BLEU-4 | METEOR | CIDEr | Acc. | Acc. | |
[…]
w/o Question | 2.56 | 9.06 | 7.65 | 0.42 | 4.97 | 12.73 | 8.44 | 7.96 |
with Question | 3.32 | 10.03 | 11.61 | 0.89 | 4.27 | 15.59 | 11.21 | 7.50 |
Substep 3-2. Once a question has been posed and the answer has been provided, the second learning phase guides the model to continue with subsequent reasoning steps.
During substep 3-1, we utilize only the part of the dataset that precedes the answers to the questions. This allows the model to focus solely on generating the lead-up to the question within the reasoning steps. For the learning in substep 3-2, the input provided to the model includes the image, instruction text, and the reasoning steps up to the response to the query. The training objective is then to output the continued reasoning steps from that point onwards.

### v2

@@ Abstract
The increasing demand for intelligent systems capable of interpreting and reasoning about visual content requires the development of large Vision-and-Language Models (VLMs) that are not only accurate but also have explicit reasoning capabilities. This paper presents a novel approach to develop a VLM with the ability to conduct explicit reasoning based on visual content and textual instructions. We introduce a system that can ask a question to acquire necessary knowledge, thereby enhancing the robustness and explicability of the reasoning process. To this end, we developed a novel dataset generated by a Large Language Model (LLM), designed to promote chain-of-thought reasoning combined with a question-asking mechanism. The dataset covers a range of tasks, from common ones like caption generation to specialized VQA tasks that require expert knowledge. Furthermore, using the dataset we created, we fine-tuned an existing VLM. This training enabled the models to generate questions and perform iterative reasoning during inference. The results demonstrated a stride toward a more robust, accurate, and interpretable VLM, capable of reasoning explicitly and seeking information proactively when confronted with ambiguous visual input.
@@ 1 Introduction
In recent years, LLMs have gained significant popularity in the field of artificial intelligence [1, 24, 31]. Building upon this, research has steped into large Vision-and-Language Models (VLMs) [11, 5, 18, 17]. Our study specifically focuses on tasks involving both vision and language modalities. Generally, VLMs learn the alignment between images and texts using large datasets sourced by web crawling. They are then fine-tuned with smaller, high-quality datasets for instruction tuning to enhance their text generation capabilities for various tasks. However, these models often struggle with hallucination, where they produce outputs not aligned with the given input. For instance, in VLMs, the models may mention objects that do not exist in the given images [15, 14].
@@
Therefore, our research aims to address these challenges by incorporating an explicit reasoning process and the ability to generate questions during reasoning. This approach is akin to the Chain-of-Thought reasoning seen in LLMs [36, 4], where models are prompted to explicitly reason out problems. However, due to VLMs’ relatively weaker text generation capabilities compared to LLMs, achieving Chain-of-Thought solely through prompting is challenging. To overcome this, we have created a novel dataset that includes the explicit reasoning process. Following mainstream methods in LLM-based dataset creation [18], we generate this data using LLMs, combining image annotations and a few manually created examples. This data includes scenarios where the model needs to generate questions when uncertain, training it to ask questions during ambiguous reasoning situations. We refer to this reasoning process as “Chain-of-Reasoning (CoR).”
In order to present the effectiveness of our approach, we fine-tune an existing VLM model on the dataset we created. Being trained on our dataset, the model acquires the ability to generate explicit reasoning steps and ask questions when uncertain, thereby improving the reliability of its inferences. Our contributions are summarized as follows:
We present an approach by incorporating an explicit reasoning process and question-generation capability into VLMs, promoting more reliable inferences.
We crafted a new dataset and utilized it for model training, setting a precedent for future VLM advancements.
@@ 2 Related Work
2.1 Large-scale Vision-and-Language Models
@@
BLIP-2 [11] and InstructBLIP [5] utilize pre-trained image encoders and text decoders, employing a Transformer model known as Q-Former as an adapter. The training of these models is typically divided into two stages. In the first stage, Q-Former learns the alignment between image and text using large-scale image-text paired datasets. The second stage involves training the Q-Former alongside the text decoder to enable the decoder to produce outputs according to given instructions. Generally, while the first training stage prioritizes the scale of the dataset, the second stage emphasizes the quality of smaller, high-quality instruction and answer datasets. This two-stage learning approach is widely adopted in the training of VLMs, as seen in models like BLIVA [9], MiniGPT (v1&2) [39, 2], and LLaVA (v1&1.5) [18, 17].
Additionally, there’s growing interest in VLMs that not only produce outputs aligned with overall image instructions but also focus on specific regions within images. This research direction is evident in models such as KOSMOS-2 [26], Shikra [3], MiniGPT v2 [2], and LLaVA v1.5 [17]. These models transform bounding box coordinates, representing input regions, into special tokens for textual representation. Another noteworthy model is GPT4RoI [38], which applies an RoIAlign based on the target region to the features outputted by the image encoder, thus obtaining features focused on the specified region.
@@ 2.3 Visual Question Generation
A related field to the concept of acquiring information through question generation is Visual Question Generation (VQG). In the early stages of VQG, the primary focus was on generating questions related to images, without any specific goal of acquiring knowledge [22, 12, 16, 35]. K-VQG [33, 34] does involve generating questions for knowledge acquisition; however, there has been no emphasis on utilizing the acquired knowledge in further reasoning steps.
Learning by Asking (LBA) is another relevant area. LBA is an approach where question generation is used to acquire data for learning purposes. For instance, LBA has been employed in tasks like VQA [21, 32, 34], caption generation [29], and scene graph completion [37]. However, our research significantly diverges from LBA, as we focus on generating questions as part of explicit reasoning during the inference process. The acquired answers are then immediately used to update and refine the reasoning quality, differentiating our approach from traditional LBA work.
@@ 3.1 Dataset Construction
In pursuit of realizing the CoR, we developed a novel dataset. This dataset was created using a combination of the specially designed prompt, manually curated examples and the text generation capabilities of LLM. We show an overview of the dataset construction process in Figure 2 and an example of the dataset in Figure 3.
The data structure of our dataset consists of the image, an instruction text of the task, a question (optional, if the task is in QA format), the answer, and the reasoning process required to derive the answer. We used different approaches to build the dataset depending on the types of annotations provided with the base dataset.
@@
(1) If the dataset had rich annotations, including captions and object detection annotations, such as labels and bounding boxes, we followed the LLaVA approach [18]. In this case, we used an LLM (GPT-4 [25]) that does not process images to build the dataset. Coupling these inputs with thoughtfully designed prompts (Listing LABEL:lst:prompt) enables the LLM to generate the appropriate reasoning steps. It’s crucial that the prompts ensure that the LLM simulates the act of observing the image directly, without mentioning any references to captions, descriptions, or bounding boxes. This approach ensures the consistency of the reasoning process when training VLMs, which do view the image directly. To achieve this, in the prompt given to the LLM, we clearly state like you must not use the word “caption”, “description”, or “bounding box” in the reasoning steps. Especially, do not mention the coordinates of each object in the image. If you want to mention the position of the object, please refer to it by relative position, such as “at the left of the image” or “in front of xx”. Furthermore, in the prompt, we instruct the LLM to suppose that the image is of size 10001000 pixels. This ensures accurate position estimations based on given textual coordinates of the bounding boxes. The prompts also emphasize succinct reasoning steps without redundancy, and the need to add uncertainty scores at the end of each reasoning step.
(2) If the dataset did not have such rich annotations, we used a model that can input images (GPT-4V) to build the dataset. We used the similar prompts as previous case, excluding the parts that required GPT to imagine image input from text information.
@@
We devised three variants of reasoning step data: without QA, with QA, and with GT. In “without QA” setting, the reasoning process to the answer is conducted without generating any questions midway. This setting aims to mimic the process a VLM undergoes when attempting to answer a question directly, without generating any questions. In this scenario, the LLM is instructed to generate reasoning steps without seeing the correct answer to the given task. We instruct the model to generate “Uncertainty score” along with each reasoning step, indicating the model’s confidence in its reasoning. This plays a crucial role in “with QA” setting described below.
@@
In contrast, the “with QA” setting incorporates the generation of questions as part of the reasoning steps. This setup is crucial for training VLMs to perform explicit reasoning while simultaneously generating questions. As the model generates the reasoning steps, it also generates an uncertainty score for each step. If the score significantly rises at any point in the “without QA” setting, it triggers the generation of a question just before the step. This mechanism helps the model to identify moments of uncertainty and seek external information, thereby simulating a more interactive and dynamic reasoning process.
@@
Lastly, the “with GT” setting involves the LLM generating reasoning steps while having access to the correct answers to the tasks (i.e., the ground-truth answers for the given questions or captions for the given images). This approach differs from the previous two as the reasoning steps invariably lead to the correct answer, serving as a ground truth for correct reasoning without question generation. This data is primarily used to train VLMs that do not involve question generation in their reasoning process.
@@
| COCO Caption | 5,857 | 5,782 | 6.27 | 8.75 | 6.20 |
| VQA v2 | 5,755 | 5,633 | 4.68 | 7.36 | 4.54 |
| OK-VQA | 5,793 | 5,792 | 4.77 | 7.35 | 4.55 |
| A-OKVQA | 5,736 | 5,718 | 4.87 | 7.43 | 4.63 |
| Visual Genome | 5,883 | 5,609 | 4.34 | 7.31 | 4.12 |
| Encyclopedic VQA | 6,521 | 6,186 | 7.00 | 9.45 | 7.00 |
| OVEN | 5,685 | 5,685 | 6.99 | 9.51 | 6.99 |
| Total | 41,230 | 39,272 | 5.58 | 8.19 | 5.46 |
@@ 3.2 Dataset Statistics
We provide a comprehensive overview of our dataset statistics in Table 1.
In our research, we based our dataset generation on seven datasets: COCO Caption [13], Visual Genome Caption [10], VQA v2 [6], OK-VQA [19], A-OKVQA [28], OVEN [8], and Encyclopedic-VQA [20].
We can categorize these datasets into three groups:
(1) Visual understanding tasks.
This group includes COCO Caption, Visual Genome Caption, and VQA v2, which are captioning tasks and VQA tasks that require the model to understand or describe the content of the image.
Usually, these tasks do not require the model to have expert knowledge that cannot be directly observed from the image.
(2) Vision + common-sense understanding tasks.
This group includes OK-VQA and A-OKVQA, which require the model to have common-sense knowledge to answer the questions.
Here, common-sense knowledge refers to general knowledge about topics such as geography, brands, and vehicles. [19]
(3) Vision + encyclopedic knowledge tasks.
This group includes Encyclopedic-VQA and OVEN, which require the model to have encyclopedic knowledge to answer the questions.
Encyclopedic knowledge refers to more in-depth and specific knowledge compared to common-sense knowledge.
This includes knowledge such as the names of animal and plant species, or details about when and by whom a building in an image was constructed [20].
We assume that the difficulty of the reasoning process increases as we move from visual understanding tasks to vision + encyclopedic knowledge tasks, as the latter requires detailed and expert knowledge to answer the questions. For each of these datasets, we generated approximately 6,000 data samples, randomly selecting from their training sets. We show examples of each category in Figure 4.
From Table 1, we can see that when generating reasoning data without questions and answers (without QA) and when not seeing ground-truth answers (with GT), the average number of reasoning steps is shorter compared to the scenario where both questions and answers are included (with QA). This is attributed to the fact that the QA generation process inherently requires more reasoning steps.
@@
The COCO Caption, Encyclopedic-VQA, and OVEN dataset tend to have longer reasoning steps compared to the others. For COCO Caption, this is likely due to the nature of the task, which involves generating captions based on the entire content of the image, requiring a more comprehensive reasoning process. The Encyclopedic-VQA and OVEN datasets often contains questions that require very detailed observations or expert knowledge to answer, leading to longer reasoning steps. In contrast, the other tasks focus on specific questions or target regions within the images, thereby requiring less information to be incorporated into the reasoning steps.
@@ 3.3 Model Architecture
Our model fundamentally builds upon the LLaVA architecture [18], consisting of an image encoder, a text decoder, and an adapter component based on two-layer MLP blocks. We show an overview of our model in Figure 5. The input image is first resized to 336336 pixels and passed through a pre-trained image encoder (e.g., CLIP-ViT [27]). The image encoder extracts the image features, which are then fed into the Adapter MLPs to transform the image features into a format that can be processed by the LLM. The instruction text is also fed into the LLM, and the model generates a text response based on the image features and instruction text.
OK-VQA | A-OKVQA | VQA v2 |
OVEN | average | |||||||
| LLaVA (original) | 2.258 | 1.424 | 2.414 | 2.332 | 2.604 | 1.606 | 1.521 | 2.023 | ||||||
| CoR w/o question | 1.621 | 1.653 | 2.224 | 2.054 | 2.524 | 1.520 | 1.582 | 1.883 | ||||||
| CoR w/ uncertainty | 1.748 | 1.748 | 2.536 | 2.408 | 2.715 | 1.902 | 1.761 | 2.117 | ||||||
| CoR w/ uncertainty, w/o knowledge | 1.775 | 1.739 | 2.526 | 2.376 | 2.717 | 1.909 | 1.839 | 2.126 | ||||||
| Ours CoR | 1.769 | 1.782 | 2.631 | 2.459 | 2.737 | 1.925 | 1.836 | 2.163 |
3.4 Training
We utilizes our dataset to make the model learn to generate reasoning steps and questions. In the training process, we utilized a pre-trained VLM, which is already instruction-tuned, and fine-tuned it on our dataset. In the fine-tuning process, we froze the parameters of the image encoder, and updated the parameters of the text decoder and the adapter MLP. Along with the image and original question, we provided a specially designed prompts, such as “Analyze the image and outline your reasoning process step by step before providing your final answer,” to let the model generate reasoning steps and questions.